Find (a) the reduced form. (b) the numerical value of π‘Œπ‘’ , and (c) the effect on the multiplier if a proportional income tax (t) is incorporated into the model. π‘Œ = 𝐢 + 𝐼, 𝐢 = 𝐢𝑂 + π‘π‘Œ 𝑑 , 𝑇 = 𝑇𝑂 + π‘‘π‘Œ, π‘Œ 𝑑 = π‘Œ βˆ’ 𝑇 where 𝐼 = 𝐼𝑂 = 30, 𝐢𝑂 = 85, 𝑏 = 0.75, 𝑑 = 0.2, and 𝑇𝑂 = 20


1
Expert's answer
2021-03-12T09:37:43-0500

(a) "T=To+tY"

"C=Co+mpc(Y-tY)"

"Y=AE=C+I+G"

"Y=Co+mpc(Y-tY)+I+G"

"Y+mpctY-mpcY=C+I+G"

"Co+MPC(Y-tY)"

"Y=\frac{1}{1-mpc(1-t)}(C+I+G)"

("Y=\frac{1}{1-mpc(1-t)}(C+I)"

(b)"Y=\frac{1}{1-mpc(1-t)}(C+I)"

"Y=\frac{1}{1-0.75(1-0.2)}(85+30)"

"Y=\frac{1}{1-0.75(1-0.2)}(85+30)"

Y=575

(c)"\frac{1}{1-mpc(1-t)}"

the multiplier increases if a proportional income tax is included in the model